### v1

# LoC-Path: Learning to Compress for Pathology Multimodal Large Language Models

Abstract
Whole Slide Image (WSI) understanding is fundamentally challenging due to its gigapixel scale and extreme sparsity of diagnostically relevant regions. Unlike human experts who primarily rely on key areas to arrive at the diagnosis, existing slide-level multimodal large language models (MLLMs) for pathology rely on heavy slide-level encoders that process thousands of tile/patch features in a brute-force manner, resulting in excessive computational cost. In this work, we revisit the WSI–language modeling paradigm and reveal that tile-level features exhibit strong global and local redundancy, whereas only a small subset of informative tiles are truly task-relevant. Motivated by these observations, we introduce an efficient MLLM framework, called LoC-Path, that replaces the expensive slide-level encoder with redundancy-reducing modules. We first design a Sparse Token Merger (STM) and an MAE-pretrained resampler to remove local redundancy and compress globally redundant tile tokens into a compact slide-level representation set. We then propose a Cross-Attention Routing Adapter (CARA) and a Token Importance Scorer (TIS) to integrate the compressed visual representation with the language model in a computation-efficient manner. Extensive experiments demonstrate that our approach achieves performance comparable to existing state-of-the-art whole slide MLLMs, but with significantly lower computational and memory burden.
1 Introduction
Digital pathology has undergone a significant revolution in recent years. Whole Slide Images (WSIs), with many exceeding $10,000\times 10,000$ pixels, are partitioned into small patches. [^1] Throughout this paper, we use tile and patch interchangeably. Powerful pretrained image encoders [32, 7, 48] are applied to these patches to extract patch-wise features. These features are then aggregated through an attention mechanism to make a whole-slide level prediction [29, 24, 55], whereas the attention will highlight a potentially small set of critical patches, imitating how pathologists make clinical decisions.
The advent of large language models (LLMs) has further changed the WSI prediction methodology. Given whole slide images, new LLM-based methods can not only make predictions such as subtype classification or survival prediction, but also generate content-rich text-based outputs, including answers to questions (VQA) or pathology reports. In particular, adapting recent multimodal large language models (MLLMs) [1, 26, 28], methods such as WSI-LLaVA [27], SlideChat [9], TCP-LLaVA [35], and ALPaCA [15] process features of all patches with a vision encoder to convert them into visual tokens, and then concatenate these visual tokens with language tokens (from a question or other textual information) as a long sequence and send it to pretrained LLMs.
(a) Compression. Mean normalized reconstruction error vs. $K$; shaded band shows interquartile range across slides. We mark the elbow and the $K$ needed to reach $50\%$ reconstruction error.
(b) Local redundancy. Fraction of tiles whose nearest neighbor exceeds cosine threshold $\tau$; IQR across slides. Values above each dot indicate the mean fraction.
(c) Pairwise similarity. Aggregated cosine histogram with vertical lines at the $\tau$ used in (b); annotations report $P(\cos>\tau)$. The inset zooms into the high‑similarity tail.
Figure 1: WSI tile tokens are redundant globally and locally. We analyze $\sim$2k TCGA WSIs (tissue tiles at $20\times$; features from CONCH‑V1 [32]). We showed that (a) we only need a small number of prototypes to reconstruct all tokens; (b) locally, tokens are similar to each other; and (c) globally, tokens are also similar.
Despite the strong prediction and generation power, these LLM-based methods all suffer from an efficiency issue. Due to the large size of a whole slide image, there can be thousands of patches, resulting in thousands of tokens being fed to an LLM. Even with a frozen LLM backbone, the extensive forward and backward passes over thousands of tokens represent a massive computational bottleneck, resulting in prolonged training durations for fine-tuning a 7B model. Furthermore, before sending these visual tokens into LLMs, it is often necessary to improve their representation quality through global reasoning, e.g., using attentional poolers to summarize slide‑level features as additional features [14] or using LongNet [13] to capture the diagnostically relevant global dependencies [52] efficiently. Although these models are capable of capturing the global context, reasoning over thousands of visual tokens creates a substantial computational and memory burden on the vision side.
To address these efficiency limitations, we initiate a quantitative investigation to answer a fundamental question: are these thousands of tokens really necessary? The answer, we find, is “no”, for two key reasons.
Reason 1: redundancy of token content. We observe that the tile-level token sequences of WSIs are highly redundant both globally and locally. We analyze the tile features of the 2,0002{,}000 TCGA [50] WSIs and show in Fig. 1: (1) a small set of $K\!\approx\!250$ prototype groups can fairly well reconstruct the entire tile feature set; and (2) tile features are highly similar to each other, as observed by comparing neighboring tiles and comparing pairs of tiles far apart. These findings confirm strong redundancy in both global content and local neighborhoods, suggesting that WSIs can be effectively represented by a compact set of latent representations rather than the full sequences of thousands of tokens.
Reason 2: irrelevance to the task. We also observe that only a limited number of tiles are truly relevant to the task, i.e., relevant to the text supervision (answers in the VQA task, and report content in the report generation task). We use the same 2,0002{,}000 TCGA WSIs and compute the cosine similarity between each tile token and the corresponding text supervision embedding. We rank tokens by their descending cosine similarity with the text supervision and compute the cumulative sum of positive similarity as a function of total token fraction. Shown in Fig. 2, the top $35.2\%$ of tile tokens account for $90\%$ of the total cosine similarity. In other words, a majority of visual tokens are irrelevant to the pathology report. This aligns with the fact that human experts only rely on key areas of WSIs for diagnosis. More analysis details are in Sec. 6.
Our contribution. Inspired by these observations, we tackle the efficiency of pathology MLLMs. Focusing on the two issues, we propose an efficient framework named LoC-Path with novel modules to (1) reduce the token redundancy, and (2) select only task-relevant tokens for the pretrained LLM. As a consequence, we reduce the number of tokens sent to LLMs from thousands per slide to hundreds, decrease inference computation by $70\%$, and reduce memory consumption by $30\%$ without sacrificing much prediction/generation power.
To reduce global redundancy of tokens, we design a lightweight resampler inspired by how pathologists inspect WSIs. The resampler scans all the tile-level tokens and then selects those crucial ones for reconstructing the whole slide with a set of learnable queries. This transforms thousands of redundant tile tokens into a compact set. To teach the resampler how to glance broadly and select distinctively, we adopt Masked Autoencoder (MAE) pre-training [18], which encourages recovery of missing tiles from partial observations. Compared to conventional slide-level encoders, our resampler is light-weight, significantly improving computational efficiency. Furthermore, to reduce local redundancy, we introduce a sliding-window Sparse Token Merger (STM) that merges highly correlated neighboring tiles before resampling, reducing sequence length while preserving representative tissue diversity.
To select task-relevant tokens, we propose a novel Token Importance Scorer (TIS) module to select tokens that are the most relevant to the task. Using query text token information, TIS ranks the projected visual tokens and selects the top-$M$ most relevant ones. Furthermore, existing LLaVA-style fusion methods use self-attention on concatenated visual and text tokens [28]. We stipulate that this is unnecessary, and propose to use cross-attention across visual and text tokens instead. Assuming $n$ visual tokens and $n$ textural tokens, the self-attention will generate $(2n)^{2}=4n^{2}$ attention values, whereas cross-attention will only generate $n^{2}$ attention values, saving computation by 75%. We inject TIS-selected visual tokens into multiple decoder layers of the LLM through Cross-Attention Routing Adapters (CARA) equipped with learnable gating mechanisms. This multi-layer routing allows richer supervision from multiple depths of the LLM, improving alignment while also reducing computation due to the aforementioned reason.
In summary, our main contributions are as follows.
  • WSI token compression. Inspired by the observed high global and local redundancy of WSI tile tokens, we propose a lightweight, MAE-pretraining-based resampler to eliminate global redundancy. We also propose an STM to eliminate local redundancy.
  • Efficient visual token selection and multimodal fusion. We introduce TIS module that selects the most task-relevant tokens, and integrates these tokens with the pretrained LLM through gated cross-attention and dynamic token routing (CARA), yielding much lower computational complexity than standard LLaVA-style fusion.
  • Efficiency with competitive performance. On different benchmarks, our LoC-Path framework achieves comparable or superior performance compared to SOTA WSI MLLMs, while significantly reducing computation and GPU memory consumption.
Figure 2: Most tokens are not task-relevant. For each slide, we sort tokens by their cosine similarity to the slide caption (expert report) and compute the cumulative share of positive similarity as a function of the token fraction. We report the mean (solid), median (dashed), and inter-quartile range (shaded) over different slides. Reaching $50\%$, $80\%$, and $90\%$ of the positive similarity mass requires $\approx 7.7\%$, $\approx 22.8\%$, and $\approx 35.2\%$ of tokens, respectively. This shows that a small subset of tokens carries most report-relevant information.
Figure 3: The overall framework architecture of LoC-Path, MAE pretraining for Resampler, and the computational illustration of STM. (a) Foreground tile tokens, $\bm{X}$, are generated by the tile-level vision encoder. STM merges them to $\bm{\tilde{X}}$. Then the resampler takes both $\bm{\tilde{X}}$ and learnable latents $\bm{Z_{0}}$ as input. The projector aligns the embedding dimension of output latent $\bm{Z}$ with LLM’s hidden dimension, $D$. The scorer selects the top $M$ latents, $\bm{\hat{Z}}$ among $L$ latents, where $M\leq L$. The CARA modules are inserted into multiple LLM decoder layers. Within each adapter, the latents perform cross-attention with the hidden state of LLM. (b) Resampler is pretrained without STM. The tile-level tokens, $\bm{X}$ are first randomly masked and then sent to context layers. Outputs, $\bm{\hat{X}}$ from context layers serves as V and K for the following cross-attention layers; a 2D positional encoding (PE) is added only to K. During pre-training, a two-layer decoder is used to reconstruct the tokens from the resampler’s output latents. (c) The diagram shows the merging process when window size $s=2$. White blocks indicate empty positions (zero tokens) in $\mathcal{S}_{g}$. The $\text{LN}_{in}$ and $\text{LN}_{out}$ are applied per tile token along the feature dimension $D_{v}$.
2 Related Work
Multimodel Large Language Models (MLLMs) for Digital Pathology. The development of multimodal large language models (MLLMs) has revolutionized computational pathology by enabling sophisticated integration of visual and textual information for histopathological analysis [25, 4]. Recent work has introduced specialized vision-language models for pathology that combine pretrained vision encoders with large language models through instruction tuning on extensive pathology datasets, achieving state-of-the-art performance on diagnostic tasks [30, 34, 23]. Advanced systems leverage contrastive learning frameworks trained on large-scale image-caption pairs to develop foundation models capable of zero-shot transfer across diverse histopathological tasks [31, 33]. Comprehensive benchmarking studies have systematically evaluated foundation models across multiple pathology datasets, revealing that pathology-specific vision-language models consistently outperform general-purpose vision models on histopathological tasks [16, 3]. Recent efforts have extended these capabilities to whole-slide image understanding, enabling multimodal reasoning at the slide level through specialized architectures that process gigapixel images with clinical context [27, 45, 36, 9, 44].
[…]
3 Methodology
A WSI is decomposed into a long sequence of tiles. For each training image-text pair, we denote the tile-level tokens from the tile-level encoder by $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{N}]\in\mathbb{R}^{N\times D_{v}}$, with 2D-grid coordinates $\bm{C}=(c_{i},r_{i})\in\mathbb{N}^{2}$, where $N$ is the number of the tile-level tokens, $D_{v}$ is the tile-level encoder’s feature dimension. $c_{i}$ and $r_{i}$ are the column and row grid coordinates with stride $1$, respectively. In MAE pre-training task, we produce $L$-length visual latents $\bm{Z}\in\mathbb{R}^{L\times D_{v}^{\prime}}$, where $D_{v}^{\prime}$ is the embedding dimension of MAE resampler. The text query tokens are $\bm{Q}_{\text{text}}$ and the answer labels are $\bm{Y}$. In single-round visual question answering (VQA) tasks, $\bm{\hat{Y}}$ is generated by an MLLM for a given $\bm{X}$ and text query $\bm{Q}_{\text{text}}$.
Overview. Our LoC-Path framework is illustrated in Fig. 3. During inference, tile-level tokens, $\bm{X}$ are merged by STM to $\bm{\tilde{X}}$. MAE-pretrained resampler compresses the merged tokens $\bm{\tilde{X}}$ to latents $\bm{Z_{0}}\in\mathbb{R}^{L\times D_{v}^{\prime}}$. The projector converts the embedding dimension and TIS selects Top-$M$ latents, $\bm{\hat{Z}}$. The selected $\bm{\hat{Z}}$ are integrated with LLM’s multiple decoder layers through CARA modules.
3.1 Compressing Slide-level Representation
To achieve a concise yet informative slide-level representation, we emulate the way human pathologists examine WSIs in our resampler design. To teach the resampler to achieve this goal, MAE pre-training is adopted: the resampler is forced to scan the whole tile-level sequence and select the tokens that are crucial for reconstruction.
Resampler for Tile-level Token Compression. To compress the tile-level tokens, we adopted the popular resampler-style design in many open-source MLLMs [1, 54, 49]. Unlike the previous designs, we break down our design into two parts: scanning and selecting to imitate human experts. A set of context layers scans the whole tile-level sequence roughly, just like the human scanning WSIs. After scanning, we design a set of learnable latent queries, each latent selects the distinct tile-level tokens from WSI. To make each latent work separately instead of looking at the same area, a Top-K attention forces each latent to look at the input token that has Top-$K$ attention scores. Then a self-attention module allows each latent to interact with each other to prevent them from looking at the same set of tokens. The details of the resampler architecture are shown in Fig. 3 (b). Concretely, given tile-level tokens $\bm{X}\in\mathbb{R}^{N\times D_{v}}$, we first apply a set of lightweight context layers to obtain $\bm{\hat{X}}=\operatorname{ContextLayers}(\bm{X})$, where $\bm{\hat{X}}\in\mathbb{R}^{N\times D_{v}}$. Then initial $L$ learnable queries $\bm{Z_{0}}\in\mathbb{R}^{L\times D_{v}^{\prime}}$ attend to $\bm{\hat{X}}$ through $R$ stacked Top-K cross-attention blocks with a LayerNorm and a residual FFN. We inject 2D positional encoding only on key: $K=(\bm{\hat{X}}+\Phi)W_{k},\ V=\bm{\hat{X}}W_{v}$. For each block: $\bm{H}=\operatorname{softmax}(\operatorname{TopK}(\operatorname{LN}(\bm{Z_{\ell-1}})K^{\top}/\sqrt{d_{h}}))V$, $\bm{Z^{\prime}}_{\ell}=\bm{Z}_{\ell-1}+\bm{H}$, $\bm{Z^{\prime\prime}_{\ell}}=\bm{Z^{\prime}}_{\ell}+\operatorname{FFN}(\operatorname{LN}(\bm{Z^{\prime}}_{\ell}))$. A self-attention layer following the cross-attention layer further decorrelates latents. $\operatorname{TopK}$ keeps the largest $K_{top}$ entries per head and per query (others set to -$\infty$ before softmax). The final output from the resampler is the latents, $\bm{Z}\in\mathbb{R}^{L,D_{v}^{{}^{\prime}}}$, where $L\ll N$.
MAE Pre-training for Redundant Tile-level Tokens. Humans only require hundreds of WSIs to learn how to find diagnoses. Sadly, it does not hold for machines. To teach the resampler with a limited number of available WSIs (around $30k$ WSIs in the TCGA dataset [50]), we adopt the Masked Auto-encoder pre-training framework. As we have shown in Fig. 1(a), the highly redundant tokens mean that the resampler can reconstruct masked tiles from visible context without overfitting to noise. Moreover, reconstruction on randomly masked indices prevents the resampler from outputting trivial averages and forces different latents to look at complementary regions. Concretely, we add a small decoder to use the latents for reconstruction, shown in Fig. 3 (b). The pre-training objective is written as: $\mathcal{L}_{\text{rec}}=\frac{1}{|\mathcal{N}|}\sum_{i\in\mathcal{N}}\|\hat{\bm{x}}_{i}-\bm{x}_{i}\|_{2}^{2}$.
To avoid degenerate average latents and to cover the slide, we add two small terms:
$\textbf{Coverage:}\quad\mathcal{L}_{\mathrm{cover}}$ $=\frac{1}{N}\sum_{i}\max\!\big(0,\;\tau-u(i)\big),$ (1)
$u(i)$ $=\max_{z}p_{z}(i),\quad$ (2)
$\textbf{Diversity:}\quad\mathcal{L}_{\mathrm{feat}}$ $=\big\|\mathrm{offdiag}(\mathrm{Cov}(\bm{Z}))\big\|_{F}^{2}.$ (3)
where $p_{z}(i)$ is attention probability of a latent code $z$ on token $i$ and $\tau$ is an adaptive-parameter, explained in detail in Sec. 8. The total pretraining loss is $\mathcal{L}_{\mathrm{MAE}}=\mathcal{L}_{\mathrm{rec}}+\lambda_{\mathrm{cover}}\mathcal{L}_{\mathrm{cover}}+\lambda_{\mathrm{feat}}\mathcal{L}_{\mathrm{feat}}$.
Sparse Token Merger. To reduce obvious local redundancy before sending the tile-level tokens to the resampler, we merge tile-level tokens using a lightweight module, STM. On a virtual 2D coordinate grid, STM uses a non-overlapping $s\times s$ window for merge. For each window $g$ with index set $\mathcal{S}_{g}$, where the number of $\mathcal{S}_{g}$ is determined by the window size: $g=u\cdot\lfloor r_{i}/s\rfloor+\lfloor c_{i}/s\rfloor$, where $u$ is a large number to prevent hash collision. For each $\mathcal{S}_{g}$, we first compute $\tilde{\bm{x}}_{i}=\mathrm{LN}_{\mathrm{in}}(\bm{x}_{i}),i\in\mathcal{S}_{g}$ and merge them within $\mathcal{S}_{g}$:
$\tilde{\bm{x}}_{g}=\mathrm{LN}_{\mathrm{out}}\!\Big(\frac{1}{\sqrt{|\mathcal{S}_{g}|}}\sum_{i\in\mathcal{S}_{g}}\tilde{\bm{x}}_{i}\Big),\quad\tilde{\bm{c}}_{g}=\big(\!\lfloor r_{i}/s\rfloor,\lfloor c_{i}/s\rfloor\!\big)$ (4)
yielding $\tilde{\bm{X}}\in\mathbb{R}^{N^{\prime}\times D_{v}}$ with $N^{\prime}\!\approx N/s^{2}$ and merged coordinates $\bm{\tilde{C}}$. For training numerical stability while we use mixed precision, we empirically divide $\sqrt{|S_{g}|}$ for each $S_{g}$ instead of $|S_{g}|$. The demo computation procedure is shown in Fig. 3 (c).
3.2 Integrating Vision with LLM Efficiently
While our resampler (Sec. 3.1) reduces tile tokens to $L$ latents, to avoid the costly fusion within LLM, we propose an efficient two-stage fusion mechanism. First, a 2-layer MLP projector aligns latents to the LLM space: $\bm{V}=\mathrm{Proj}(\bm{Z})\in\mathbb{R}^{L\times D}$, where $D$ is the embedding dimension of LLM. Second, a lightweight Token Importance Scorer (TIS) dynamically selects the Top-$M$ query-relevant latents ($M\leq L$). These are fed into our Cross-Attention Routing Adapters (CARA), which efficiently integrate them with the LLM at a cost of $O(TM)$, where $T$ is the length of the text tokens.
[…]
Cross-Attention Routing Adapter (CARA). We integrate the selected latents $\bm{V}_{\mathcal{I}}$ into the LLM. Inspired by Flamingo [1], we insert CARA into a small set of decoder layers $\mathcal{L}$. Given hidden states $\bm{H}^{(\ell)}$ at layer $\ell$, CARA performs gated cross-attention: $\tilde{\bm{H}}^{(\ell)}=\bm{H}^{(\ell)}+g_{\ell}\cdot\mathrm{Attn}\!\Big(\mathrm{LN}(\bm{H}^{(\ell)}),\bm{V}_{\mathcal{I}},\bm{V}_{\mathcal{I}}\Big)$ where $g_{\ell}$ is a learnable gate (initialized small) to stabilize early training. Then, the modality integration loss, with the standard auto-regressive loss $\mathcal{L}_{\mathrm{LM}}$ between the model output $\bm{\hat{Y}}$ and text label $\bm{Y}$ can be written as
$\mathcal{L}=\mathcal{L}_{\mathrm{LM}}+\lambda_{\mathrm{TIS}}\mathcal{L}_{\mathrm{TIS}}+\lambda_{\mathrm{rank}}\mathcal{L}_{\mathrm{rank}}.$ (5)
4 Experiment
4.1 Experimental Settings
Dataset. We use three distinct datasets to evaluate our design choices. The WSI-Bench dataset, introduced in WSI-LLaVA [27], is the primary source for training the resampler and for subsequent modality fusion with LLMs. It has over $9k$ WSIs and over $170k$ VQA pairs. For zero-shot testing, we assess the trained models on the WSI-Bench test set, as well as on Slide-Bench from SlideChat [9] and selected data from Path-Bench [46]. Notably, Slide-Bench contains BCNB WSIs [51], which differ from the TCGA WSIs [50] in WSI-Bench, making it an effective out-of-distribution (OOD) test set. As both WSI- and Path-Bench are based on the TCGA dataset, we remove any overlapping WSIs between WSI- and Path-Bench training sets. After this adjustment, we retain 400 WSIs and 780 visual question-answer pairs from the Path-Bench dataset. For the pre-processing, we first apply DSMIL [24] and CLAM [29] to identify tissue patches at $20\times$ magnification and remove background regions. By using the CONCH V1 [32], we then extract tile-level features from these patches. The tile size is $224\times 224$. More details are provided in Sec. 8.
Experimental Details. The resampler is first pre-trained using MAE for $100k$ steps. After pre-training the resampler, we fuse the modality with the LLM in another two training stages. We use Qwen2.5-7B-Instruct [47] as the LLM to balance training cost and performance. The adapters are inserted to decoder layers $\{1,3,5,7\}$ in LLM. In the first fusion stage, we freeze the resampler and LLM, but with STM on and only train the projector, TIS, and CARA with $M=96$. During this stage, we train the model using the report data in WSI-Bench [27]. In the second fusion stage, we use all the VQA pairs and unfreeze the latents and the last layer of the resampler. More details are in Sec. 8.
Baselines. To verify the effectiveness of our LoC-Path framework, we compare it with several end-to-end MLLMs, including GPT-4o [21], Quilt-LLaVA [41], WSI-LLaVA [27], and SlideChat [9]. The input tokens for SlideChat [9] are $20480$. We resize the WSI thumbnails to $1024\times 1024$ for Quilt-LLaVA [41]. To compare with other token reduction methods, we compare Token Merge (ToMe) [56, 5], DivPrune [2], ACMIL [55], and the adaptive average pooling method. We replace our STM and the resampler with these token reduction methods, but leave the projector, TIS, and CARA components unchanged. We set the merged result token length to be $256$ with $M=96$. For ToMe, we use the balanced implementation from AIM [56], and we iteratively do tile-level token merge until the token length reaches $256$. For ACMIL [55], we follow the attention head design and use the attention score to select $256$ tile-level tokens. For DivPrune [2], we prune token length at different rates to reach the same $256$ token length. We also train the token reduction baselines in two fusion stages.
Evaluation Metrics.
For the visual question answering (VQA) tasks, we follow the approach of WSI-LLaVA [27] and categorize the tasks into four types: morphological analysis, diagnosis, treatment planning, and report generation. For open-ended questions, we evaluate performance using BLEU [39], WSI-Precision, and WSI-Relevance, while accuracy is used for closed-ended tasks. WSI-Precision and WSI-Relevance are LLM-based evaluation metrics, proposed in WSI-LLaVA [27], with GPT-4o [21] serving as the evaluator. To assess the effectiveness of our resampler and the quality of the slide-level latent representations, we use AUC [17] and Accuracy [43] as the primary metrics for classification tasks. For the feature quality assessment from the resampler, we assess the latent attention coverage at $\tau{=}10^{-4}$ with $\mathbb{E}\big[\frac{1}{N}\sum_{i}\mathbf{1}\{\max_{z}p_{z}(i)>\tau\}\big]$, similar definition in Equ. 3 as well as the pairwise cosine similarity (diversity) between each latent code.
4.2 Results
Table 1: Comprehensive comparison of our resampler with $256$ latents and GigaPath’s slide-level encoder on classification performance, feature quality, and computational cost. Feature quality is measured on $2k$ WSIs sampled from the TCGA dataset based on the tumor type. Computational cost (including peak memory and TFLOPS) is measured during inference with a fixed input of $60k$ tokens.
Model Subtype Vital Analysis Feature Quality (mean) Computational Cost
AUC $\uparrow$ Acc $\uparrow$ AUC $\uparrow$ Acc $\uparrow$ Diversity (std) $\downarrow$ Coverage $\uparrow$ Mem (MB) $\downarrow$ TFLOPS $\downarrow$
Ours (no regularizer) 0.741 0.672 0.714 0.845 0.110 (0.462) 0.812 2254.5 1.713
Ours (MAE) 0.763 0.734 0.696 0.845 0.346 (0.056) 0.894 2254.5 1.713
Slide-level Encoder [52] 0.727 0.672 0.725 0.854 0.382 (0.197) 4022.2 10.340
Figure 4: Coverage and visualization of latent attention preferences. The left bottom panel shows coverage: a token $i$ is marked covered by any latent $z$ if $\max_{z}p_{z,i}>\tau$; coverage = fraction of covered tokens ($\tau=10^{-4}$). 80% coverage means the resampler attends to 80% of token positions on the slide. $z_{0}$ denotes the spatial locations of the top-$64$ input visual tokens receiving the highest attention score from the resampler’s latent $z_{0}$. Orange points mark selected tokens; the blue background shows all token positions.
Compressed Slide-level Representation from the Resampler. To verify the design of our resampler, we first compare it with the slide-level encoder on two classification tasks: sub-typing and vital analysis of BRCA tumor type. As shown in Tab. 1, our light-weight resampler with 256 latents achieves comparable AUC on both classification tasks compared to GigaPath [52]. To achieve such performance, our resampler only requires half the memory and $\approx 1/6$ of FLOPs compared to the slide-level encoder. For the feature quality, we mainly check resampler’s latent attention coverage and latent diversity, qualitatively shown in Fig. 4. In Tab. 1, latents from the resampler achieve similar diversity as GigaPath while maintaining low std across WSIs. For the coverage, the latents can cover over $80\%$ input tokens. Two regularization terms further improve the attention coverage and keep the latents diverse. The details about the classification head implementation, and more comparisons of other pre-training methods are in Sec. 9.
Figure 5: Query-aware visual tokens selection. Each heatmap shows the aggregated attention of the Top-$M$ TIS‑selected latents for a single WSI under two different queries. The brighter regions correspond to visual tokens that the TIS module deems most relevant to the query, illustrating how the model focuses on distinct morphological cues depending on the question.
Figure 6: The quality comparison between our model and the reference answer. The same color indicates the matched claims between the model’s response and the reference.
The Overall Performance against Baselines. In Tab. 2, our framework outperforms all token reduction baselines and achieves comparable performance with other end-to-end MLLMs, especially WSI-LLaVA (ours 0.739 vs 0.754). In Tab. 3, although our method on Slide-Bench is a bit lower than DivPrune and Avg. Pooling, our method achieves the highest WSI-P (0.384) and WSI-R (0.473) in the report generation task, even when comparing to WSI-LLaVA [27]. Although other conventional NLU metrics are lower than WSI-LLaVA [27], higher WSI-P and WSI-R have been shown to be more accurate and clinically relevant than conventional NLU metrics in pathological contexts [27]. Our model is the most computationally efficient in Tab. 2 due to the design of TIS and CARA modules. The inference TFLOPs of our model (L=128/M=96) are $81.91\%$ lower than all the general baselines compared, and inference memory is also $38.9\%$ lower. The qualitative comparison to the reference answer is shown in Fig. 6. More analysis are in Sec. 12.
Fusion Style Comparison. We compare fusion styles in Tab. 4 (b), fixing the resampler output to $L=256$ latents and routing $m=96$ tokens. Our CARA+TIS fusion method achieves the highest performance on WSI-Bench, Avg (0.738). The LLaVA-style fusion achieves the second highest WSI-Bench average (0.733). The standard Flamingo-style (no projector and only attention projection heads used in cross-attention adapters) baseline is slightly lower on the WSI-Bench average. Adding the TIS to the Flamingo-style adapters recovers performance. This comparison indicates that using projector to integrate vision features with LLM improves the overall performance like LLaVA-style and our CARA+TIS improves the performance compared to standard LLaVA-style fusion.
Table 2: Detailed performance and efficiency comparison on the WSI-Bench test set, categorized by task type. WSI-P and WSI-R denote WSI-Precision and WSI-Relevance, respectively, evaluated by GPT-4o [21]. Acc denotes Accuracy. Inference memory and TFLOPs are compared for all General $\downarrow$ (lower is better). Values are per single forward pass, reporting peak reserved memory usage (GB) and TFLOPs. We compare only the LLM generation cost (visual inputs are pre-extracted for SlideChat [9] and WSI-LLaVA [27]). LLMs loaded in 16-bit with SDPA attention and KV cache on; ours merges LoRA weights [19]. Percentages for Ours are reductions vs. the mean of General.
Model Max Visual Token Number Memory (GB) $\downarrow$ TFLOPs $\downarrow$ Morphological Analysis Diagnosis Treatment Planning Average $\uparrow$
Open Close Open Close Open Close
WSI-P $\uparrow$ WSI-R $\uparrow$ Acc $\uparrow$ WSI-P $\uparrow$ WSI-R $\uparrow$ Acc $\uparrow$ WSI-P $\uparrow$ WSI-R $\uparrow$ Acc $\uparrow$
General:
GPT-4o [21] - - - 0.220 0.204 0.471 0.472 0.457 0.530 0.496 0.841 0.875 0.507
Quilt-LLaVA [41] 576 12.620 12.798 0.448 0.447 0.947 0.586 0.604 0.849 0.788 0.816 1.000 0.721
SlideChat [9] 20480 11.583 26.767 0.269 0.281 0.870 0.319 0.385 0.767 0.654 0.577 0.333 0.495
WSI-LLaVA [27] 576 13.619 10.491 0.488 0.610 0.951 0.610 0.612 0.863 0.810 0.845 1.000 0.754
Token Reduction:
ToMe (AIM) [56] 256 - - 0.535 0.593 0.942 0.528 0.556 0.842 0.733 0.779 1.000 0.723
DivPrune [2] 256 - - 0.527 0.581 0.921 0.491 0.515 0.820 0.518 0.627 0.854 0.650
ACMIL [55] 256 - - 0.535 0.593 0.943 0.528 0.556 0.842 0.733 0.779 1.000 0.723
Avg. Pooling 256 - - 0.532 0.593 0.940 0.545 0.574 0.845 0.666 0.766 0.958 0.713
Ours (L=256/M=96) 256 7.709 3.388 0.564 0.626 0.943 0.570 0.598 0.860 0.734 0.771 0.979 0.738
Ours (L=128/M=96) 128 7.709  (-38.9%) 3.019  (-81.91%) 0.567 0.620 0.931 0.582 0.613 0.849 0.733 0.769 1.000 0.739
Table 3: Combined results on WSI-Bench (report generation) and zero-shot generalization (Slide-Bench and Path-Bench subset). Rows under General are general-purpose MLLMs; rows under Token Reduction apply token compression.
Model WSI-Bench: Report Generation Zero-shot Generalization
BLEU-1 $\uparrow$ BLEU-2 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ WSI-P $\uparrow$ WSI-R $\uparrow$ Slide-Bench Acc $\uparrow$ Path-Bench Acc $\uparrow$
General:
GPT-4o [21] 0.202 0.069 0.132 0.167 0.067 0.138 0.414 0.793
Quilt-LLaVA [41] 0.474 0.351 0.475 0.460 0.324 0.333 0.163 0.380
SlideChat [9] 0.439 0.107 0.161 0.137 0.123 0.168 0.546 0.847
WSI-LLaVA [27] 0.480 0.358 0.490 0.465 0.380 0.429 0.553 0.690
Token Reduction:
ToMe (AIM) [56] 0.586 0.309 0.384 0.418 0.377 0.469 0.508 0.820
DivPrune [2] 0.556 0.286 0.375 0.417 0.335 0.419 0.544 0.737
ACMIL [55] 0.572 0.294 0.374 0.401 0.334 0.422 0.508 0.821
Avg. Pooling 0.569 0.296 0.372 0.406 0.320 0.423 0.545 0.837
Ours (L=256/M=96) 0.579 0.312 0.390 0.426 0.369 0.462 0.519 0.813
Ours (L=128/M=96) 0.583 0.313 0.392 0.425 0.384 0.473 0.519 0.847
4.3 Ablation Study
Table 4: Ablation results on latent length sweep and different modality fusion strategies. WSI-Bench (except report generation) reports the average score ($\uparrow$). For the ablation on fusion style, $L$ and $M$ are fixed at $256$ and $96$, respectively. Full per-task results are provided in Sec. 11.
(a) Latent length sweep
$L$ / $M$ WSI-Bench Avg
128 / 96 0.739
256 / 96 0.738
256 / 144 0.729
512 / 384 0.741
1024 / 576 0.731
(b) Fusion style
Fusion WSI-Bench Avg
LLaVA 0.733
Flamingo 0.726
Flamingo + TIS 0.729
Ours w.o. TIS 0.736
Ours 0.738
Impact of the Latent Length in Resampler. We focus on the MLLM setting and vary the resampler’s latent length $L$ while keeping the Top-$M$ selected tokens active to capture language‑side complexity. On WSI‑Bench (except report generation), a moderate latent pool works best: L=512L{=}512 with M=384M{=}384 attains the highest average (0.741), compared to others in Tab. 4 (a). The trend indicates that increasing $L$ initially improves retrieval of rare but relevant regions for the router, whereas overly large $L$ introduces redundancy and dilutes attention. If we keep the $M$ while increasing $L$, the performance decreases a little bit.
Table 5: STM impact on vision cost (TFLOPs, Memory) vs. final MLLM performance (WSI-Bench Avg).
Setting $s$ TFLOPs $\downarrow$ Mem (MB) $\downarrow$ WSI-Bench Avg $\uparrow$
No Merge 0 1.713 2254.5 0.735
Merge 2x 2 0.440 704.5 0.738
Merge 4x 4 0.121 607.1 0.737
Impact of STM. On the vision side, using STM translates into large measured savings: with the same 60k‑token input, adding STM with s=4s{=}4 saves around $14\times$ TFLOPs and $3.7\times$ memory reduction relative to the resampler‑only setting, without much performance decrease, shown in Tab. 5. Also, the averaged score with s=2s{=}2 slightly exceeds the no‑merge setting, with per‑task changes mostly within $\pm$0.02. Taken together with the large compute savings above, STM offers a favorable accuracy–efficiency trade‑off.
Impact of TIS. The effectiveness of TIS is shown in two comparisons, in Tab. 4 (b). First, comparing our full model to an ablation without it, the inclusion of TIS provides a clear performance gain of 0.002. Second, to isolate its effect further, we added TIS to a standard “Flamingo” style fusion. The result shows the same positive trend: “Flamingo + TIS” outperforms the standard “Flamingo” baseline by 0.003 margin. In both contexts, the TIS consistently improves the model’s average score. This demonstrates that dynamically routing only the most salient visual information, as guided by the TIS, qualitatively shown in Fig. 5, is effective to help the LLM focus and improves performance.
5 Discussion and Conclusion
In this work, we addressed the computational inefficiency of applying multimodal large language models (MLLMs) to gigapixel-scale WSIs by leveraging their inherent data redundancy. We quantitatively demonstrate that this redundancy, both local and global, aligns with the sparse diagnostic process of human pathologists. We propose the LoC-Path framework, featuring two redundancy-aware modules: a Sparse Token Merger (STM) to prune local redundancy and an MAE-pretrained Resampler to compress the global tile sequence. To efficiently fuse this representation, our novel CARA+TIS mechanism avoids quadratic complexity by restricting LLM attention to only the Top-$M$ most salient visual latents, mimicking expert focus. Our extensive experiments demonstrate that our method achieves diagnostic performance comparable to that of state-of-the-art MLLMs while substantially reducing computational requirements (FLOPs) and memory demands. Crucially, beyond efficiency, the framework offers inherent interpretability through intermediate signals—such as STM merging grids, resampler attention maps, and TIS rankings—providing intuitive views of how the model allocates capacity across diagnostically relevant regions.
Limitations. Despite these gains, LoC-Path currently serves as an assisting tool rather than a standalone clinical system. First, the model relies on fixed pre-trained encoders; systematic biases or domain shifts (e.g., rare tumor subtypes) in the feature extractor inevitably propagate to the LLM. Second, while effective for global patterns, our compression prioritizes coverage over worst-case preservation, potentially under-representing subtle but decision-critical lesions. Consequently, as shown in our qualitative analysis in Sec. 12, the model struggles with fine-grained precision in complex tasks: it may hallucinate specific attributes (e.g., exact Nottingham scores) in report generation or offer generic rather than patient-specific regimens in treatment planning. Finally, current automatic metrics (e.g., WSI-Precision) may overestimate reliability by failing to penalize clinically dangerous hallucinations.
Future Research Directions. To bridge the gap between a reference tool and a hospital-deployable system, future work will focus on several key areas. First, we aim to extend the resampler’s MAE pre-training to handle diverse data granularities simultaneously, unifying patch-level and slide-level pathology data to learn more robust, multi-scale representations. Second, regarding modality fusion, we will explore broader possibilities for the CARA+TIS module, investigating diverse fusion modes to better integrate compressed visual signals with language logic. Furthermore, to improve report generation fidelity and treatment planning details, we plan to scale up training data and switch to more powerful LLM backbones, while integrating structured supervision—such as clinical guidelines and genomics—to reduce hallucinations. Finally, rigorous human evaluation and uncertainty estimation are essential to validate LoC-Path’s safety for real-world clinical deployment.
Supplementary Material
6 Redundancy and Task Relevance Analysis
We quantify how much of a WSI’s tile–token stream is redundant (globally and locally), and how many tokens are actually task–relevant to the textual supervision. Our analyses operate on pre–extracted tile embeddings and are computed per slide and then summarized across slides by mean and inter–quartile range (IQR).
[…]
Global Compression Redundancy. To quantitatively show that the tile-level tokens are compressible, we design the following experiments: For $K\in\{8,16,\dots,256\}$ we learn $K$ prototypes $\{ce_{k}\}$ (K‑means). We assign each $x_{i}$ to $ce(x_{i})$ and compute the normalized reconstruction error:
$\mathrm{nMSE}(K)=\frac{\frac{1}{N}\sum_{i=1}^{N}\lVert x_{i}-ce(x_{i})\rVert_{2}^{2}}{\frac{1}{N}\sum_{i=1}^{N}\lVert x_{i}-\bar{x}\rVert_{2}^{2}}.$
Per slide, we obtain a curve $K\mapsto\mathrm{nMSE}(K)$; across slides, we show the mean and IQR, and annotate the elbow and the $K$ achieving $\approx$50% error reduction. As a result, in Fig. 1(a), we show that a small $K$ prototypes are enough to reconstruct the slide fairly well, indicating potential for compression.
Local Redundancy. To assess the similarity between one tile and its neighbors, we find its nearest spatial neighbors $x_{nn(i)}$ and evaluate cosine similarity. For thresholds $\tau\in\{0.90,0.92,0.94,0.96,0.98\}$ we report
$r_{\tau}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left[\cos\!\bigl(x_{i},x_{nn(i)}\bigr)>\tau\right].$
, where $N=16$ in our experiment. We aggregate $r_{\tau}$ over slides by mean and IQR. A large fraction of tiles exceed high cosine thresholds with a neighbor, which indicates strong local redundancy as we show in Fig. 1(b).
Global Pairwise Similarity. Is high similarity common beyond nearest neighbors? To answer this question, we sample up to $50k$ unordered tile pairs $(x_{i},x_{j})$ per slide and build a cosine histogram aggregated over slides. The right-skewed histogram in Fig. 1(c) demonstrates the high global redundancy.
Task Relevance Sparsity. To assess how many tiles are actually supervised by the report (caption), we study the relations between tile-level tokens and the global report embedding. Let $t$ be the slide’s text (report) embedding. We follow the method in WSI-LLaVA [27] to clean the report, leaving the morphological and diagnosis information. For each tile, we compute $s_{i}=\cos(x_{i},t)$ and then sort tiles by $s_{i}$ in descending order and form the cumulative share of the positive similarity mass,
$F(p)=\frac{\sum_{i=1}^{\lfloor pN\rfloor}\max(s_{i},0)}{\sum_{i=1}^{N}\max(s_{i},0)},\qquad p\in[0,1].$
[…]
7 Pairwise Ranking Loss Details
As mentioned in Sec. 3.2, we add a pairwise ranking loss ($\mathcal{L}_{\mathrm{rank}}$) to the TIS training objective to stabilize learning and enforce order consistency.
To compute this loss, we first sort the Top-$M$ indices $I$ according to the teacher distribution $\bm{q}$, yielding $I_{\text{sorted}}$. We select $n=\min(M/2,N_{\text{pairs}})$ pairs, where $N_{\text{pairs}}$ is a hyperparameter (e.g., 8). We define the ”high” set $I_{\mathrm{hi}}=I_{\text{sorted}}[:n]$ (the top $n$ indices) and the ”low” set $I_{\mathrm{lo}}=I_{\text{sorted}}[M//2:M//2+n]$ (the first $n$ indices from the bottom half).
We form a set of $n$ matched pairs $\mathcal{P}=\{(i_{k},j_{k})\mid i_{k}\in I_{\mathrm{hi}},j_{k}\in I_{\mathrm{lo}},k=1..n\}$, where pairs are matched by their rank (i.e., 1st-highest $i_{1}$ is paired with 1st-lowest $j_{1}$). The loss uses a margin $\gamma$ and weights $w_{ij}$ that emphasize pairs with larger teacher gaps (i.e., pairs the teacher is more confident about):
$w_{ij}$ $=\frac{\max(q_{i}-q_{j},\,0)}{\frac{1}{|\mathcal{P}|}\sum_{(i^{\prime},j^{\prime})\in\mathcal{P}}\max(q_{i^{\prime}}-q_{j^{\prime}},\,0)},$ (6)
$\mathcal{L}_{\mathrm{rank}}$ $=\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}w_{ij}\,\big[\,\gamma-(s_{i}-s_{j})\,\big]_{+},$ (7)
where $\cdot]_{+}=\max(\cdot,0)$, and $s_{i}$ denotes the raw (unnormalized) scores from the TIS.
8 Implementation and Training Details
Implementation Details. The resampler has 2 LongNet [13] context layers and 2 cross-attention layers with self-attention layers as we discussed in Sec. 3.2. The segment set used for dilation attention in two LongNet [13] layers is [1024, 2024, 4096]. The stride sets are [1, 2, 4] and [4, 8, 16]. Moreover, we enable the relative positional encoding within LongNet’s dilation attention to help context layers better capture the short-to-mid range context [13]. For the following cross-attention layers, we set $K=128$ for the first cross-attention layer and $K=64$ for the second cross-attention layer. The embedding dimension of the resampler is 512, and the attention head is 8 for both context layers and cross-attention layers. We use 2D Fourier positional encoding for the resampler’s cross-attention key values. Doing so can help us deal with the filtered (missing) coordinates in the tile-level sequence and save memory, compared to 2D sinusoidal positional encoding.
Classification Task Conversion Details. In feature quality comparison experiments, we demonstrate the feature quality of the latents from the resampler. We use two classification tasks using BRCA tumor types. The first vital analysis classification tasks mark survival as 1 and not as 0; the sub-typing tasks define the first stage as 1 while other labels as 0, with unknown stages removed. The classification task head is a simple attention-based classification head from ABMIL [22].
Training Details. The resampler is pre-trained for $100\mathrm{k}$ iterations on CONCH-V1 [32] WSI features with a mini-batch size of $2$ per GPU and gradient accumulation $4$ on $4$ RTX A6000 Ada GPUs (effective global batch size $32$). We randomly mask $75\%$ of the input tokens and optimize the MAE objective $\mathcal{L}_{\mathrm{MAE}}$ introduced in Sec. 3.1, with $\lambda_{\mathrm{cover}}=5\times 10^{-4}$ for the coverage term $\mathcal{L}_{\mathrm{cover}}$ and $\lambda_{\mathrm{feat}}=1\times 10^{-3}$ for the diversity term $\mathcal{L}_{\mathrm{feat}}$. We use AdamW with an initial learning rate $1\times 10^{-4}$, a $3\%$ warm-up phase, and a cosine learning-rate schedule. During this pre-training we gradually anneal the Top-$K$ in the resampler cross-attention from $(128,64)$ to $(64,32)$ following a cosine schedule.
We then train the first fusion stage for $3$ epochs with a global batch size of $512$ (batch size $2$ per GPU and gradient accumulation $64$ on $4$ GPUs) and an initial learning rate $1\times 10^{-4}$. In this stage we freeze both the Qwen2.5-7B-Instruct LLM and the resampler, and only train the visual projector together with the CARA+TIS modules on top of the LLM. The training objective is the multimodal loss as defined in Sec. 3.2, where we set $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$ for the self-distillation (TIS) loss $\mathcal{L}_{\mathrm{TIS}}$ and the ranking loss $\mathcal{L}_{\mathrm{rank}}$.
In the second fusion stage, we unfreeze the resampler’s last cross-attention layer and its latent in-projection, and insert LoRA modules (rank $128$, $\alpha=64$) into all linear layers of the LLM. We then train all unfrozen components on all VQA pairs in WSI-Bench for $2$ epochs with a global batch size of $128$ and an initial learning rate $2\times 10^{-5}$, again using the same multimodal fusion loss with $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$. Across all training stages, we use Xtuner [10] with DeepSpeed ZeRO-2 [40] and FlashAttention 2 [11] for efficient training; FA2 is applied to the LLM self-attention but not to the cross-attention layers. A cosine learning-rate scheduler is used in all stages, and all experiments are conducted on $4$ RTX A6000 Ada GPUs.
Training Optimization. The pre-training requires about $10$ hours. The first fusion stage takes about $1$ hour. The second fusion stage requires about $8$ hours to $9$ hours.
During development of the core training pipeline, we observed that the primary bottleneck was not GPU computational capability, but the overhead of transferring data to the GPUs. In digital pathology, tile-level features are commonly stored in .pt or .h5 formats. While these formats are not a problem for training on one or two GPUs, they incur substantial I/O overhead at larger scales, causing the data loader to spend a significant amount of time waiting for data. To mitigate this issue, we convert all tile-level features and their corresponding coordinates into .npy files separately, which supports more efficient memory mapping and lightweight loading. Furthermore, we store the features in float16 format while keeping coordinates in float32 format, to further reduce both storage and I/O costs. Since we only use mixed-precision training throughout all stages of our pipeline, using features in float16 format is not an issue.
With file format conversion, we are able to load long sequences with very low I/O costs. Unlike the conventional tile-level data loader in digital pathology [9, 51], loading one data point in a single batch, we are able to load multiple tile-level sequences in a single batch simultaneously. To deal with sequences with different lengths, we pack these sequences into the same length (we pad sequences with different lengths to the same $60k$ lengths and load padded sequences with their corresponding padding masks). As a result, we can increase the batch size to more than one (e.g 4) on each GPU. Doing so allows us to reduce the data loading rounds if we set the max training epochs to be $2$.
With these optimizations, the training time of the second fusion stage is reduced from more than a day to less than 9 hours on our system. We also observe 2-3 times training time reduction on other stages.
9 Resampler Pretraining Method Comparison
Table 6: Comparison of resampler pre-training strategies. We report AUC and Accuracy for BRCA tumor subtyping and vital analysis tasks. Latent quality is evaluated via Mean Pairwise Cosine and Union Coverage, calculated on $2k$ random samples from the WSI-Bench training set.
Encoder Pre-training Subtype Vital Analysis Feature Quality (mean)
AUC $\uparrow$ Acc $\uparrow$ AUC $\uparrow$ Acc $\uparrow$ Diversity (std) $\downarrow$ Coverage $\uparrow$
CONCH V1 [32] MIL-NCE [37] 0.692 0.719 0.754 0.857 0.999 ($3.0\text{e-}4$) 0.065
LLM-loss [28] 0.724 0.688 0.759 0.882 0.841 (0.049) 1.000
MoCoV3 [8] 0.705 0.594 0.554 0.863 0.999 ($4.1\text{e-}5$) 0.288
Ours (no regularizer) 0.741 0.672 0.714 0.845 0.110 (0.462) 0.812
Ours (MAE) 0.763 0.734 0.696 0.845 0.346 (0.056) 0.894
Prov-GigaPath [52] MAE (Slide-level) 0.727 0.672 0.725 0.854 0.382 (0.197)
We have already shown the MAE-pretrained resampler’s output latent (feature) quality in Sec. 4.2 and it has achieved comparable quality compared to Gigapath’s slide-level encoder [52]. Here, we compare our pre-training strategy against other pre-training strategies. These methods include MIL-NCE pre-training method [37] using randomly sampled tile-level tokens and their caption (report) pairs, pre-training only using LLM’s auto-regressive loss (LLaVA-style fusion), and pre-training using MoCoV3 [8]. For MIL-NCE pre-training [37], on the vision side, during training, we randomly sample over 500 tile-level tokens; on the language side, we use the global embedding of cleaned reports in WSI-Bench [27]. For the text encoder, we use CONCH-V1 [32]’s corresponding text encoder. For the end-to-end LLM-loss pre-training, we follow the LLaVA-like first fusion stage pre-training by adding a two-layer projector between LLM and resampler [28]. For the MoCoV3 [8] pre-training method, we first create two different views by sub-sampling the same tile-level token sequence for two resampler instances. In Tab. 6, the MAE-pretrained resampler’s internal cross-attention exhibits a good diversity score and high attention coverage, as we have shown in Sec. 4.2. Furthermore, its output latents demonstrate strong orthogonality (average cosine similarity 0.35). This indicates that it genuinely samples broadly across the whole slide rather than simply averaging. It also achieves an AUC of 0.763 on the BRCA subtype task.
The pre-pretraining methods using text-supervision, including MIL-NCE [37] and end-to-end LLM-loss [28], achieve much higher downstream AUC and Acc on the vital analysis task compared to those that do not. This suggests that the text supervision does help the resampler create the link between the input to the final diagnosis prediction. The MIL-NCE pretrained resampler has very low attention coverage, while producing very similar output latents (very poor diversity). This indicates that this pre-training method only encourages the resampler to pay attention to those tile-level tokens with relevant text supervision. It is problematic when it comes to more rare cases, such as answering questions regarding locally specific morphological features. The LLM-loss-pretrained resampler exhibits high attention coverage while having similar output latents (very poor diversity). Combining these two pieces of evidence indicates that this resampler just computes the average for the input tile-level tokens instead of picking the important ones that can represent the whole tile-level sequence. The MoCoV3 [8] pretrained resampler has low attention coverage with very similar output latents. Thus, the MAE pretraining method actually achieves our design goal: the resampler first glances the whole sequence (high attention coverage), and then each latent code looks at the distinct place on the tile-level tokens rather than computing an average.
Table 7: Ablation of resampler architecture. We evaluate the impact of latent length ($L$), cross-attention depth, and context layer type on BRCA subtype and vital analysis tasks (AUC/Accuracy). Unless otherwise specified, the default configuration uses $L=256$, MAE pre-training, and regularization terms.
Resampler Config. Sub-typing Vital Analysis
AUC $\uparrow$ Acc $\uparrow$ AUC $\uparrow$ Acc $\uparrow$
Ours 0.763 0.734 0.696 0.845
w/o Reg. Loss 0.741 0.672 0.714 0.845
Num. Latents:
     L=128L{=}128 0.800 0.688 0.720 0.826
     L=512L{=}512 0.739 0.672 0.662 0.789
     L=1024L{=}1024 0.760 0.719 0.687 0.839
Cross-Attn Depth:
     3 Layers 0.764 0.688 0.750 0.845
Context Layer:
     Depth-wise Conv. 0.741 0.656 0.674 0.826
Model Morphological Analysis Diagnosis Treatment Planning Average $\uparrow$
WSI-P WSI-R Acc WSI-P WSI-R Acc WSI-P WSI-R Acc
Fusion Ablation
LLaVA 0.556 0.619 0.933 0.558 0.594 0.852 0.722 0.759 1.000 0.733
Flamingo 0.551 0.605 0.934 0.537 0.572 0.835 0.729 0.771 1.000 0.726
Flamingo + TIS 0.552 0.615 0.944 0.536 0.566 0.834 0.756 0.782 0.979 0.729
Gating Ablation (7B)
TIS Off 0.561 0.614 0.935 0.571 0.603 0.845 0.716 0.783 1.000 0.736
Adapter Insertion Layers (7B)
Adapter First 0.557 0.611 0.931 0.563 0.601 0.847 0.711 0.770 0.979 0.730
Adapter Layers 1/8/15/22 0.558 0.609 0.933 0.553 0.591 0.847 0.734 0.777 1.000 0.734
Token Merging Ablation
No Merge 0.556 0.614 0.934 0.553 0.597 0.860 0.725 0.772 1.000 0.735
Merge Size 3 0.557 0.608 0.935 0.566 0.594 0.840 0.707 0.792 0.979 0.731
Merge Size 4 0.557 0.614 0.937 0.573 0.612 0.850 0.717 0.776 1.000 0.737
Latent + Router Top-$M$
L=256L{=}256, M=144M{=}144 0.555 0.608 0.940 0.549 0.587 0.827 0.718 0.774 1.000 0.729
L=512L{=}512, M=384M{=}384 0.558 0.610 0.940 0.567 0.609 0.854 0.743 0.791 1.000 0.741
L=1024L{=}1024, M=576M{=}576 0.556 0.613 0.940 0.545 0.580 0.842 0.738 0.784 0.979 0.731
Table 8: Detailed ablation results on WSI-Bench (7B LoC-Path). This table expands upon the study in Sec. 4.3 of the main paper. Metrics include open-ended WSI-Precision (WSI-P), WSI-Relevance (WSI-R), and closed-ended Accuracy (Acc) across morphological, diagnostic, and treatment planning tasks. Best results within each ablation group are highlighted in bold.
Variant Morphological Analysis Diagnosis Treatment Planning Avg. $\uparrow$
WSI-P WSI-R Acc WSI-P WSI-R Acc WSI-P WSI-R Acc
LLM Scale (CONCH-V1, STM+Resampler)
LoC-Path-7B (L=256L{=}256, M=96M{=}96) 0.564 0.626 0.943 0.570 0.598 0.860 0.734 0.771 0.979 0.738
LoC-Path-14B (L=256L{=}256, M=96M{=}96) 0.557 0.612 0.937 0.569 0.600 0.849 0.759 0.816 1.000 0.744
Adapter Insertion
Adapter at Layer 1 0.557 0.611 0.931 0.563 0.601 0.847 0.711 0.770 0.979 0.730
Adapters at Layers 1/8/15/22 0.558 0.609 0.933 0.553 0.591 0.847 0.734 0.777 1.000 0.734
CARA Gating and Ranking Loss
w/o CARA gating ($g_{\ell}{=}1$) 0.559 0.613 0.928 0.552 0.595 0.852 0.741 0.784 1.000 0.736
w/o TIS ranking loss 0.555 0.612 0.934 0.564 0.596 0.832 0.721 0.755 0.979 0.728
Tile-level Encoder
DINOv2, $D_{v}=768$, $3\times$ CA 0.454 0.484 0.898 0.418 0.472 0.809 0.665 0.635 0.917 0.639
CONCH-V15, $D_{v}=768$, $3\times$ CA 0.559 0.619 0.938 0.562 0.610 0.865 0.744 0.805 1.000 0.745
Top-$K$ Attention in Resampler
Dense resampler (w/o Top-$K$) 0.511 0.579 0.916 0.470 0.505 0.739 0.621 0.696 0.854 0.655
Table 9: Extended ablation studies. We analyze the impact of LLM scale, adapter placement, CARA gating/ranking, tile-level encoders, and resampler attention patterns. Default configuration: 7B LLM, CONCH-V1 encoder, and a 512-dim resampler with two cross-attention layers. Scores represent WSI-Bench performance on non-report-generation tasks. For the LLM scale ablation, the 14B variant uses Qwen2.5-14B-Instruct as the LLM. For adapter insertion, we vary the decoder layers where CARA is added. For CARA gating, we disable the learnable gate by fixing $g_{\ell}{=}1$. For tile-level encoder ablations, we use CONCH-V1.5 and DINOv2, each with a 768-d resampler and three cross-attention layers. For the resampler ablation, the dense variant replaces the default Top-$K$ cross-attention with dense attention.
Figure 7: Visualization of MAE resampler reconstruction. Left: Original tile-level feature map (visualized by feature norm). Middle: Masked input (75% mask ratio); black regions indicate missing tokens. Right: Features reconstructed by the MAE-pretrained resampler (256 latents). Despite heavy corruption, the model successfully recovers the global tissue layout and feature magnitude, demonstrating the efficacy of the compressed latent representation.
10 Complexity Analysis of CARA Impact
[…]
During decoding, LLaVA’s per‑step attention scales with the whole prefix $(T+L)$, yielding a total cost $\mathcal{O}(nA(T+L)+nA^{2})$ for $A$ generated tokens. In our model, self‑attention is unchanged while cross‑attention is linear in $M$ and independent of $A$ on the key/value side, giving $\boxed{\mathcal{O}(n(AT+A^{2})+n_{\times}AM)}$ . Memory usage follows the same trend: LLaVA’s KV cache grows as $\mathcal{O}(nd(T+L+A))$, whereas ours is $\mathcal{O}(nd(T+A))+\mathcal{O}(n_{\times}Md)$ with a constant visual term. The gains are most pronounced when $L$ is large (e.g., WSI), since we replace the $\mathcal{O}(L)$ dependence with $\mathcal{O}(M)$ both in prefill compute and in decoding’s per‑step cost.
This theoretical analysis aligns with our empirical results in Tab. 2 of the main paper, where we observe an over $70\%$ reduction in FLOPs and $\approx 38\%$ savings in peak memory compared to standard LLaVA baselines [28].
11 More Ablation Study
Ablation of resampler architecture. Tab. 7 summarizes several design choices for the MAE-pretrained resampler on the BRCA subtype and vital-analysis classification tasks. Adding the coverage and diversity regularizers improves the subtype AUC from $0.741$ to $0.763$ and also stabilizes training, while keeping vital-analysis performance unchanged. Varying the number of latents shows that moderate compression works best: L=128L{=}128 achieves the highest subtype AUC but slightly worse vital-analysis accuracy, whereas very long latent sets (L=512L{=}512 or $1024$) underperform on both tasks. Using three cross-attention layers improves both subtype and vital-analysis AUCs compared to the default two-layer design, but at a higher computational cost; our default strikes a balance between accuracy and efficiency. Finally, replacing the LongNet-style context layers with depth-wise convolutions noticeably hurts performance, confirming the importance of long-range context modeling before resampling.
Detailed WSI-Bench Ablations. Tab. 8 reports full per-task results on WSI-Bench (except report generation task) corresponding to the ablations in Sec. 4.3 of the main paper.
Impact of Top-$K$ Attention in the Resampler. We further study whether the sparse Top-$K$ cross-attention used in the resampler is necessary. To this end, we replace the Top-$K$ operator with dense cross-attention while keeping all other components fixed (7B LLM, CONCH-V1 input tile-level tokens [32], STM, L=256L{=}256 latents and M=96M{=}96 selected tokens). As reported in Tab. 9, this dense resampler substantially hurts the WSI-Bench performance: the overall average drops from $0.738$ (LoC-Path-7B with Top-$K$) to $0.655$. The degradation is consistent across all three sub-type tasks. These results confirm that enforcing sparsity via Top-$K$ attention is crucial for learning diverse, coverage-aware latents that support strong slide-level reasoning.
Impact of TIS Ranking Loss. The new experiments in Tab. 9 isolate the effect of the ranking loss used to train the TIS. Removing the pairwise ranking loss while keeping the distillation loss leads to a noticeable drop in WSI-Bench average from $0.739$ to $0.728$. Most of the degradation comes from diagnosis and treatment-planning questions: the closed-form diagnosis accuracy decreases from $0.849$ to $0.832$, and treatment-planning obtains lower WSI-P/WSI-R. These results suggest that the ranking loss helps TIS produce a more stable and well-ordered importance distribution over latents, which in turn improves cross-layer consistency of the routed tokens.
Impact of CARA Insertion Layers. To study where to insert CARA modules inside LLM, we compare placing only one CARA module in the LLM’s layer index 1 (“Adapter at Layer 1”) with distributing them across layers at index $\{1,8,15,22\}$, as reported in Tab. 8. Both variants use the same visual configuration (L=256L{=}256, M=96M{=}96), but the distributed design achieves a higher WSI-Bench average ($0.734$ vs. $0.730$). This indicates that injecting visual information at multiple depths allows the LLM to refine its reasoning with fresh slide-level evidence, rather than relying solely on early-layer fusion. However, when it compares to our default settings (inserting the CARA at LLM decoder layer index $\{1,3,5,7\}$), distributed insertion gives slightly decreased performance. This might be caused by the fact that the deep decoder layers of LLM do not pay attention to the visual information, unlike the early layers.
Impact of the Gating Mechanism. Our CARA modules include a learnable scalar gate $g_{\ell}$ that controls how much visual information is injected at each decoder layer. In Tab. 8, turning off the gate (“TIS Off”) slightly reduces the WSI-Bench average and harms open-ended WSI-P/WSI-R on diagnosis, even though all visual latents are still available to the LLM. The experiments in Tab. 9 show an even clearer trend: forcing $g_{\ell}{=}1$ for all layers reduces the WSI-Bench average from $0.744$ to $0.736$, with noticeable drops in WSI-Precision and WSI-Relevance for both diagnosis and treatment planning. Together, these results confirm that letting the model softly gate visual information is more effective than always fusing all latents, and it also saves computation by avoiding unnecessary cross-attention when the gate is small.
Impact of Tile-level Encoders. We primarily use CONCH-V1 [32] features (512 feature dimension) as tile-level inputs, but our framework is compatible with other tile-level vision encoders. Tab. 6 already shows that the MAE-pretrained resampler produces high-quality latents from CONCH-V1 features [32], matching the slide-level encoder upper bound on classification tasks. To further test robustness, we swap the tile encoder to CONCH-V15 [14], whose feature dimension is $768$. To accommodate this higher dimension, we increase the resampler hidden size to $768$ and use three cross-attention layers. As summarized in Tab. 9, the CONCH-V15[14] variant improves all three WSI-Bench tasks and achieves the best overall average ($0.745$), especially boosting WSI-Relevance for diagnosis and treatment planning. This suggests that LoC-Path can directly benefit from advances in pathology foundation models without changing the overall MLLM architecture. We also swap the tile-level encoder to a Dino V2 [38] variant without feature alignment pre-training, which yields worse overall performance on WSI-Bench tasks. This suggests that the feature space alignment training is essential to bridging the vision encoders and the pre-trained LLMs.
Impact of LLM Scale. Finally, we scale the language backbone from Qwen2.5-7B-Instruct (used in the main paper) to Qwen2.5-14B while keeping the visual stack and routing modules fixed [47]. On WSI-Bench, the 7B model in the main paper reaches an average score of $0.739$, whereas the 14B variant in Tab. 9 improves this to $0.744$. The gains are consistent across morphological analysis, diagnosis, and treatment planning, indicating that our token-compression pipeline already supplies a strong visual representation even for a relatively small LLM. Scaling up the language model yields additional accuracy, but is less noticeable compared to the gain by switching to a better tile-level encoder.
Figure 8: STM merging behavior on a TCGA WSI. From left to right: (i) input WSI thumbnail, (ii) original tile grid with per-tile feature norms, and (iii) merged grid after applying STM with a local window (e.g., s=2s{=}2). STM aggressively reduces the number of tiles while largely preserving the overall tissue footprint and coarse morphology, illustrating how local redundancy can be removed before resampling.
12 More Qualitative Results
MAE Resampler Reconstruction. To better illustrate what the MAE-pretrained resampler learns, Fig. 7 visualizes an example WSI before and after reconstruction. The left panel shows the original tile-level feature map, with feature norms encoded as a heatmap over the 2D grid. The middle panel shows the masked input used during pretraining, where $75\%$ of the tiles are randomly removed. The right panel presents the reconstruction from our resampler decoder. Despite the heavy corruption, the model successfully recovers the global tissue layout and the overall magnitude of tile features, supporting our assumption that a small set of latents can faithfully encode the slide-level representation. This qualitative result complements the quantitative comparison with the slide-level encoder in Tab. 7.
Latent Diversity and Coverage. To further analyze the behavior of the resampler latents, Fig. 10 visualizes the Top-$K$ attention maps of multiple latents on the same WSI as we have shown in Fig. 4. In Fig. 10, different latents focus on distinct tissue regions and morphological patterns. At the same time, they cover most of the slide, matching the high coverage reported in Sec. 4.2. Compared with the aggregate statistics in Fig. 10 and Tab. 7, this qualitative view confirms that the resampler indeed “glances” over the whole WSI and allocates different latents to complementary visual contexts rather than redundantly attending to the same locations.
STM Merged Grid Visualization. Fig. 8 provides a visual explanation of how STM removes local redundancy. In Fig. 8, the right panel shows the merged grid after applying STM with a local window (s=2s{=}2). Although the number of tokens is significantly reduced, the merged tokens still form a dense coverage of the tissue region and preserve coarse anatomical structures. This visualization is consistent with the quantitative results in Tab. 7 and Tab. 5 in the main paper, where STM does not hurt accuracy while dramatically reducing FLOPs and memory on the vision side.
Task-level Successes and Failure Modes. Fig. 9 illustrates LoC-Path’s performance across four tasks in WSI-Bench. Regarding successes (top row), the model accurately identifies dominant tumor types and key morphological patterns, such as invasive ductal carcinoma, infiltrative growth, and nuclear pleomorphism. This alignment with expert reports confirms that our compression preserves clinically salient regions. However, failure modes exist (bottom row). In treatment planning, the model tends to offer generic guideline-based regimens while under-emphasizing case-specific constraints like biomarkers. Similarly, in report generation, while correctly identifying tumor type and grade, it occasionally hallucinates fine-grained attributes (e.g., Nottingham scores or microcalcifications). Overall, LoC-Path reliably captures coarse diagnostic intent but lacks precision for granular decision-critical details, which also aligns with the quantitative results before.
Figure 9: Qualitative success and failure cases across WSI-Bench tasks. Top row: Successful cases in morphological analysis and diagnosis, showing high overlap with expert descriptions. Bottom row: Typical failure modes in treatment planning and report generation. While the model captures global intent, it may miss targeted options or hallucinate details. (Red highlights indicate hallucinations or errors; Green highlights indicate correct key findings.)
Figure 10: Diversity of resampler latents via Top-$K$ attention. Each panel corresponds to one latent from the MAE-pretrained resampler. This resampler has 256 output latents. Orange dots mark the spatial locations of the Top-64 input tokens receiving the highest attention from that latent; the blue background shows all foreground token positions. Different latents specialize to distinct regions and structures while jointly covering the whole slide, qualitatively confirming the high coverage and diversity reported in Sec. 4.2 and Tab. 7.

### v2

@@
 A WSI is decomposed into a long sequence of tiles.
-For each training image-text pair, we denote the tile-level tokens from the tile-level encoder by $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{N}]\in\mathbb{R}^{N\times D_{v}}$, with 2D-grid coordinates $\bm{C}=(c_{i},r_{i})\in\mathbb{N}^{2}$, where $N$ is the number of the tile-level tokens, $D_{v}$ is the tile-level encoder’s feature dimension.
-$c_{i}$ and $r_{i}$ are the column and row grid coordinates with stride $1$, respectively.
+For each training image-text pair, we denote the tile-level tokens from the tile-level encoder by $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{N}]\in\mathbb{R}^{N\times D_{v}}$, with 2D-grid coordinates $\bm{C}=(c_{i},r_{i})\in\mathbb{N}^{2}$, where NN is the number of the tile-level tokens, $D_{v}$ is the tile-level encoder’s feature dimension.
+$c_{i}$ and $r_{i}$ are the column and row grid coordinates with stride 11, respectively.
 In MAE pre-training task, we produce $L$-length visual latents $\bm{Z}\in\mathbb{R}^{L\times D_{v}^{\prime}}$, where $D_{v}^{\prime}$ is the embedding dimension of MAE resampler.
@@
 For each tile, we compute $s_{i}=\cos(x_{i},t)$ and then sort tiles by $s_{i}$ in descending order and form the cumulative share of the positive similarity mass,
-$F(p)=\frac{\sum_{i=1}^{\lfloor pN\rfloor}\max(s_{i},0)}{\sum_{i=1}^{N}\max(s_{i},0)},\qquad p\in[0,1].$
+$F(p)=\frac{\sum_{i=1}^{\lfloor pN\rfloor}\max(s_{i},0)}{\sum_{i=1}^{N}\max(s_{i},0)},\quad p\in[0,1].$
 […]
@@
 As mentioned in Sec. 3.2, we add a pairwise ranking loss ($\mathcal{L}_{\mathrm{rank}}$) to the TIS training objective to stabilize learning and enforce order consistency.
-To compute this loss, we first sort the Top-$M$ indices $I$ according to the teacher distribution $\bm{q}$, yielding $I_{\text{sorted}}$.
+To compute this loss, we first sort the Top-$M$ indices II according to the teacher distribution $\bm{q}$, yielding $I_{\text{sorted}}$.
 We select $n=\min(M/2,N_{\text{pairs}})$ pairs, where $N_{\text{pairs}}$ is a hyperparameter (e.g., 8).
@@
 Training Details.
-The resampler is pre-trained for $100\mathrm{k}$ iterations on CONCH-V1 [32] WSI features with a mini-batch size of $2$ per GPU and gradient accumulation $4$ on $4$ RTX A6000 Ada GPUs (effective global batch size $32$).
+The resampler is pre-trained for $100\mathrm{k}$ iterations on CONCH-V1 [32] WSI features with a mini-batch size of 22 per GPU and gradient accumulation 44 on 44 RTX A6000 Ada GPUs (effective global batch size $32$).
 We randomly mask $75\%$ of the input tokens and optimize the MAE objective $\mathcal{L}_{\mathrm{MAE}}$ introduced in Sec.
@@
 During this pre-training we gradually anneal the Top-$K$ in the resampler cross-attention from $(128,64)$ to $(64,32)$ following a cosine schedule.
-We then train the first fusion stage for $3$ epochs with a global batch size of $512$ (batch size $2$ per GPU and gradient accumulation $64$ on $4$ GPUs) and an initial learning rate $1\times 10^{-4}$.
+We then train the first fusion stage for 33 epochs with a global batch size of $512$ (batch size 22 per GPU and gradient accumulation $64$ on 44 GPUs) and an initial learning rate $1\times 10^{-4}$.
 In this stage we freeze both the Qwen2.5-7B-Instruct LLM and the resampler, and only train the visual projector together with the CARA+TIS modules on top of the LLM.
@@
 In the second fusion stage, we unfreeze the resampler’s last cross-attention layer and its latent in-projection, and insert LoRA modules (rank $128$, $\alpha=64$) into all linear layers of the LLM.
-We then train all unfrozen components on all VQA pairs in WSI-Bench for $2$ epochs with a global batch size of $128$ and an initial learning rate $2\times 10^{-5}$, again using the same multimodal fusion loss with $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$.
+We then train all unfrozen components on all VQA pairs in WSI-Bench for 22 epochs with a global batch size of $128$ and an initial learning rate $2\times 10^{-5}$, again using the same multimodal fusion loss with $\lambda_{\mathrm{TIS}}=\lambda_{\mathrm{rank}}=0.02$.
 Across all training stages, we use Xtuner [10] with DeepSpeed ZeRO-2 [40] and FlashAttention 2 [11] for efficient training; FA2 is applied to the LLM self-attention but not to the cross-attention layers.
-A cosine learning-rate scheduler is used in all stages, and all experiments are conducted on $4$ RTX A6000 Ada GPUs.
-Training Optimization. The pre-training requires about $10$ hours. The first fusion stage takes about $1$ hour. The second fusion stage requires about $8$ hours to $9$ hours.
+A cosine learning-rate scheduler is used in all stages, and all experiments are conducted on 44 RTX A6000 Ada GPUs.
+Training Optimization. The pre-training requires about $10$ hours. The first fusion stage takes about 11 hour. The second fusion stage requires about 88 hours to 99 hours.
 During development of the core training pipeline, we observed that the primary bottleneck was not GPU computational capability, but the overhead of transferring data to the GPUs.
@@
 As a result, we can increase the batch size to more than one (e.g 4) on each GPU.
-Doing so allows us to reduce the data loading rounds if we set the max training epochs to be $2$.
+Doing so allows us to reduce the data loading rounds if we set the max training epochs to be 22.
 With these optimizations, the training time of the second fusion stage is reduced from more than a day to less than 9 hours on our system.
@@
 […]
-During decoding, LLaVA’s per‑step attention scales with the whole prefix $(T+L)$, yielding a total cost $\mathcal{O}(nA(T+L)+nA^{2})$ for $A$ generated tokens.
-In our model, self‑attention is unchanged while cross‑attention is linear in $M$ and independent of $A$ on the key/value side, giving $\boxed{\mathcal{O}(n(AT+A^{2})+n_{\times}AM)}$ .
+During decoding, LLaVA’s per‑step attention scales with the whole prefix $(T+L)$, yielding a total cost $\mathcal{O}(nA(T+L)+nA^{2})$ for AA generated tokens.
+In our model, self‑attention is unchanged while cross‑attention is linear in $M$ and independent of AA on the key/value side, giving $\boxed{\mathcal{O}(n(AT+A^{2})+n_{\times}AM)}$ .
 Memory usage follows the same trend: LLaVA’s KV cache grows as $\mathcal{O}(nd(T+L+A))$, whereas ours is $\mathcal{O}(nd(T+A))+\mathcal{O}(n_{\times}Md)$ with a constant visual term.